OpenStudy (anonymous):

the hypotenuse of a 45 degree-45degree-90 degree triangle is 60 ft long. how long is each leg?

OpenStudy (anonymous):

divide by the square root of two

OpenStudy (anonymous):



OpenStudy (anonymous):

since both angles are equal, both legs are equal if you call the length of the leg \(a\) then pythagoras tells you \[a^2+a^2=h^2\] \[2a^2=h^2\] \[a\sqrt2=h\] in other words the hypotenuse is the leg times the square root of two

OpenStudy (anonymous):

since your hypotenuse is \(60\) you divide by \(\sqrt2\) to get the length of the sides, i.e. they are both \(\frac{60}{\sqrt2}\)

OpenStudy (anonymous):

k whats that equal

OpenStudy (anonymous):

i dont have that as a choice